Assume the current price of maize K70 per 100 kg and the short run cost function, where Q represents bags of maize per year is:
TC = 800 + 16Q - Q2
a) What is the profit maximizing output?
1
Expert's answer
2016-03-09T08:36:42-0500
The conditions seems inconsistent with the reality: that is, the Total Cost function becomes negative when Q becomes more than 37:

TC (0) = 800 + 16 * 0 - (0 * 0) = 800
TC (38) = 800 + 16 * 38 - (38 * 38) = -36
...
TC (100) = 800 + 16 * 100 - (100 * 100) = - 7600

which means that the cost become "revenue", in fact. Formally, based on the conditions, the answer would be infinity, which is obviously impossible.